OpenStudy (unklerhaukus):

Solve the following differential equation \[y′′−2y′−8y=4\]

OpenStudy (anonymous):

Put y=t+k choose k such that 4 vanishes

OpenStudy (anonymous):

y'=t' y''=t'' t''-2t'-8(t-k)=4 t''-2t'-8t=4+8k Therefore k=-1/2 it becomes t''-2t'-8t=0

OpenStudy (unklerhaukus):

\[y''-2y'-8y=4\]\[\downarrow\]\[t''-2t'-8t=0\] \[(t'+2)(t'-4)=0\] \[t'=-2,4\]?

OpenStudy (unklerhaukus):

\[t=-2x,\qquad4x\]?

OpenStudy (unklerhaukus):

+c

OpenStudy (anonymous):

no dude these are ODEs We write it as \[(D^2-2D-8)t=0\] where d is the linear differential operator [(D-4)(D-2)]t=0 solutions to this equation are \[t=c_1e^{4x}+c_2e^{-2x}\]

OpenStudy (unklerhaukus):

\[y-k=c_1e^{4x}+c_2e^{-2x}\]\[y=c_1e^{4x}+c_2e^{-2x}+k\]

OpenStudy (anonymous):

u get it right